OpenStudy (anonymous):

Use standard trig identities to show that cos(x+(pi/2))=-sin(x). For each step indicate which identity you used.

OpenStudy (anonymous):

cos(x+(pi/2))=cos(x)cos(pi/2)(+-)sin(x)sin(pi/2) cos(pi/2)=0 sin(pi/2)=1

OpenStudy (anonymous):

more commonly just the - in between.

OpenStudy (campbell_st):

using the sum of angles theorem \[\cos(x + \pi/2) = \cos(x)\times \cos(\pi/2) - \sin(x) \times \sin(\pi/2)\] cos(pi/2) = 0 and sin(pi/2) = 1 = 0 - sin(x) x 1 = -sin(x)